Decide all values of b in the following equations that will
give one or more real number solutions.
1. 3x^2+bx-3=0
2. 5x^2+bx+1=0
3. -3x^2+bx-3=0
4. Write a rule for judging if an equation has solutions by looking at it in standard form.

Answer by checkley71(8403)

3x^2+bx-3=0
3x^2-8x-3=0
(3x+1)(x-3)=0
or
3x^2-3=0
(3x-3)(x+1)=0
-----------------------------------------
5x^2+bx+1=0
5x^2+6x+1=0
(5x+1)(x+1)=0
or
5x^2-6x+1=0
(5x-1)(x-1)=0
--------------------------------------
I'LL LEAVE THE REST FOR YOU TO SOLVE HAVING GIVEN YOU A COUPLE OF SOLUTIONS.